OpenStudy (anonymous):

The dimensions of a rectangular prism are the same as a Pythagorean triangle with the legs of 12 and 16. The volume of the prism is? Answers: 192 768 3,840 5,860

OpenStudy (anonymous):

The pythagorean triple is 12, 16 and 20 Hence the volume is 12 * 16 * 20 = 3840

OpenStudy (anonymous):

\[16^2−12^2= b^2\]

OpenStudy (anonymous):

Oh that makes since now. Thank you both!!